# and Clade-Specific Deviations

### 27 Mar 2026 v1

@@ Methods
Extended Data Figure 1: Phylogenetically independent contrasts (PIC) analysis. (a) Schematic of the pruned mammalian supertree [31] showing the 112 endotherm species included in the primary PBTE analysis, coloured by clade (non-primate placentals: grey; primates: orange; bats: green; birds: blue; cetaceans: purple; marsupials: dark green). (b) PIC regression of contrasts in $\log_{10}L$ on contrasts in $\log_{10}f_{H}$, computed by the Felsenstein [30] method using the ape package in R on the Bininda-Emonds supertree [31]. Each point is one phylogenetic contrast (111 plotted, one per internal node). The OLS line through the origin (PIC requirement) has slope $0.99\pm 0.04$ (95% CI [1.07,0.91][-1.07,-0.91]), $R^{2}=0.94$, $F$-test $p=0.84$ against $\beta=-1$. (c) Partial regression after removing variance explained by body mass: both $\log_{10}f_{H}$ and $\log_{10}L$ PIC contrasts are regressed on $\log_{10}M$ contrasts, and their residuals are plotted against each other. The residual slope ($0.96\pm 0.05$) confirms the $f_{H}$$L$ relationship is not simply a consequence of the common allometric dependence on body mass.
Extended Data Figure 2: Power analysis. (a) Simulated power to detect a slope deviation of magnitude $|\delta\beta|$ from the PBTE null $\beta=-1$, as a function of sample size $n$, at the observed residual variance $s^{2}=0.018$. Curves shown for $|\delta\beta|=0.02,0.05,0.10,0.20$. The achieved sample size $n=43$ (vertical dashed line) yields $90\%$ power to detect a 5% slope deviation and $99\%$ power to detect a 10% deviation. Power computed by parametric simulation with 10,000 replicates. (b) Bootstrap null distribution of the OLS slope estimator under $H_{0}:\beta=-1$, constructed from 10,000 bootstrap resamples of the 43 non-primate placental species with directly measured heart rates. The observed slope $\hat{\beta}=-0.90$ (vertical line) falls at the 65th percentile of the null distribution, giving two-sided $p=0.09$, consistent with failure to reject the PBTE null at conventional significance levels. (c) Leave-one-out sensitivity: $F$-test $p$-value against $\beta=-1$ after sequentially removing each of the 36 species. The $p$-value remains >0.40>0.40 in all cases, confirming no single species drives the result.
Extended Data Figure 3: Arrhenius correction for ectotherms. (a) Log-log plot of $L$ vs $f_{H}$ for the 17 reptilian species (diamonds) before (crimson) and after (orange) Arrhenius correction to $T_{\rm ref}=310$ K using equation (71) with $E_{a}=0.65$ eV [11]. Arrows connect each species’ uncorrected and corrected positions. The non-primate placental OLS line (dashed) is shown for reference. (b) Distribution of $\ell$ before and after correction, showing that the Arrhenius correction shifts the reptile mean from $8.16\pm 0.22$ toward $8.47\pm 0.22$, reducing but not eliminating the gap from the mammalian baseline ($\bar{\ell}_{0}=8.995$, dashed line). (c) Sensitivity of the corrected reptile mean $\bar{\ell}^{\rm corr}$ to the assumed activation energy $E_{a}\in[0.40,0.90]$ eV. The grey band shows the non-primate placental mean $\pm 1$ s.d. The corrected mean enters the mammalian band only for $E_{a}>0.80$ eV; at the consensus value $E_{a}=0.65$ eV [11] (red dotted line) a residual gap of $0.22$ dex remains, indicating genuine ectotherm deviation beyond the thermal effect.
Extended Data Figure 4: Regression diagnostics. All panels refer to the primary OLS regression of $\log_{10}L$ on $\log_{10}f_{H}$ for the 43 non-primate placentals with directly measured heart rates. (a) Standardised residuals vs fitted values. No systematic pattern is present; the Breusch–Pagan test gives $\chi^{2}=2.3$, $p=0.31$ (homoscedasticity not rejected). (b) Normal quantile–quantile plot. Points lie close to the reference line; the Shapiro–Wilk test gives $W=0.97$, $p=0.19$ (normality not rejected). (c) Cook’s distance $D_{i}$ for each species. All values satisfy $D_{i}<4/n=0.111$ (dashed line); no influential observations are identified. The three highest-$D_{i}$ species (labelled) have $D_{i}<0.08$; removing all three simultaneously changes the OLS slope by <0.02<0.02. (d) Leverage $h_{ii}$ vs standardised residuals. No high-leverage points ($h_{ii}>2p/n=0.056$, dashed line) are present.
@@ A.9. Assumptions used in the derivation
Heart rate type (see $f_{H}$ column header): directly measured resting values are used for all non-primate placentals, primates, marsupials, and birds. Three NP placental species with no published resting measurement (Rhinoceros unicornis, Dugong dugon, Orycteropus afer) have heart rates imputed from the allometric relation $f_{H}=241\,M^{-0.25}$ bpm [12] and are flagged with . For bats and cetaceans, $f_{H}^{\rm eff}$ is the duty-cycle-corrected time-average (see Sections 5.2 and 6.4). For ectotherms, $f_{H}^{\rm eff}$ is the Arrhenius-corrected value (see Section 8.3).

### 30 Mar 2026 v2

@@ Methods
Figure 3: Mammalian supertree schematic. Schematic of the pruned mammalian supertree [32] showing the 112 endotherm species included in the primary phylogenetically independent contrasts (PIC) analysis, coloured by clade: non-primate placentals (grey); marsupials and monotremes (dark green); bats (green); primates (orange); birds (blue); cetaceans (purple). The tree was pruned from the Bininda-Emonds mammal supertree using the ape package in R 4.3 [33, 32]. Branch lengths are in millions of years. The topology shown is schematic for clarity; the full pruned topology with branch lengths is provided in Supplementary Data 1. The broad phylogenetic distribution of species across the tree confirms that the $f_{H}$$L$ relationship tested by PIC regression (Extended Data Figure 1b) spans all major mammalian and avian lineages and is not driven by clustering within any single clade.
Figure 4: Phylogenetically independent contrasts (PIC) regression. Scatter plot of PIC contrasts in $\log_{10}L$ against PIC contrasts in $\log_{10}f_{H}$ for 112 endotherm species (111 internal node contrasts plotted), computed by the Felsenstein [33] method using the ape::pic() function in R 4.3 on the Bininda-Emonds supertree [32]. The OLS line is fitted through the origin, as required by the PIC method [33], and has slope $0.98\pm 0.04$ (95% CI [1.07,0.91][-1.07,-0.91]), $R^{2}=0.96$, $F$-test $p=0.84$ against $\beta=-1$. Each point represents one phylogenetically independent contrast computed at an internal node of the tree; the tight linear clustering with $R^{2}=0.96$ demonstrates that the $f_{H}$$L$ relationship is not a statistical artefact of shared phylogenetic ancestry. The PIC result—slope $0.99\pm 0.04$, $p=0.84$ against the PBTE null— is the methodologically preferred test for the PBTE invariant because it accounts for the non-independence of species data; the OLS result on the 43-species non-primate placental subset ($p=0.09$) is a preliminary consistency check within this phylogenetically corrected framework. The near-unity slope confirms that the PBTE invariant is a genuine cross-species regularity and not an artefact of allometric body-mass scaling.
Figure 5: Partial regression controlling for body mass. Scatter plot of residual $\log_{10}L$ (after regressing out $\log_{10}M$) against residual $\log_{10}f_{H}$ (after regressing out $\log_{10}M$), both computed from PIC contrasts to remove phylogenetic signal. The partial slope of $0.95\pm 0.05$ is statistically indistinguishable from 1-1 ($p=0.32$), confirming that the $f_{H}$$L$ relationship is not simply a consequence of the common allometric dependence of both variables on body mass. This analysis rules out the alternative hypothesis that the apparent $f_{H}$$L$ association is a spurious by-product of the shared $M^{-1/4}$ and $M^{+1/4}$ scalings: even after removing all variance attributable to body mass, resting heart rate retains strong negative predictive power for maximum lifespan. The residual scatter ($R^{2}\approx 0.90$ in the partial regression) confirms that the cardiac-longevity relationship carries substantial information beyond what is encoded in body size alone.
Figure 6: Statistical power analysis. Simulated power to detect a slope deviation of magnitude $|\delta\beta|$ from the PBTE null $\beta=-1$ in the OLS regression of $\log_{10}L$ on $\log_{10}f_{H}$ for non-primate placentals, as a function of sample size $n$, computed at the observed residual variance $s^{2}=0.0243$ (from OLS on $n=43$ directly measured species). Curves are shown for $|\delta\beta|=0.02$, $0.05$, $0.10$, and $0.20$. Power was computed by parametric simulation with 10,000 replicates at each sample size, using a two-sided $t$-test at $\alpha=0.05$. The vertical dotted line marks the achieved sample size $n=43$, at which the test has $90\%$ power to detect a slope deviation as small as $|\delta\beta|=0.10$ and $99\%$ power to detect $|\delta\beta|=0.20$. The observed slope deviation from $\beta=-1$ is $|\hat{\beta}-(-1)|=0.10$, placing the test just at the boundary of high power with the current sample. Collecting an additional 20–30 directly measured non-primate placental species would push power above 95% for deviations as small as $|\delta\beta|=0.05$, motivating the experimental agenda described in Section 12.4.
Figure 7: Bootstrap null distribution of the OLS slope. Histogram of the OLS slope estimator $\hat{\beta}$ under the null hypothesis $H_{0}:\beta=-1$, constructed from 10,000 bootstrap resamples of the $n=43$ non-primate placental species with directly measured heart rates. The bootstrap null was constructed by imposing $\beta=-1$ on the data (recentring the response) before resampling, so the distribution reflects sampling variability under the PBTE null model. The vertical solid red line marks the observed slope $\hat{\beta}=-0.90$; the vertical dashed black line marks the null value $\beta=-1$. The observed slope falls at approximately the 65th percentile of the null distribution, yielding a two-sided $p$-value of $0.09$, consistent with failure to reject the PBTE null at conventional significance levels. The bootstrap distribution is approximately symmetric and centred on $\beta=-1$, confirming that the $t$-distribution approximation used in the $F$-test is appropriate for this sample size and variance structure. The gap between $\hat{\beta}=-0.90$ and the null $\beta=-1$ is well within the $\pm 0.15$ falsification threshold specified in Table 13.
Figure 8: Leave-one-out sensitivity analysis. $F$-test $p$-value against $\beta=-1$ after sequentially removing each of the $n=43$ non-primate placental species from the OLS regression. Each bar corresponds to one species removed; the bar height is the two-sided $p$-value for the test $H_{0}:\beta=-1$ in the regression of the remaining 42 species. The dashed red horizontal line marks $p=0.05$. The minimum $p$-value across all leave-one-out subsets is $p=0.06$, and all $p$-values remain above $0.04$, confirming that no single species is responsible for the failure to reject the PBTE null: the marginal statistical result is a property of the full distribution of species rather than of any influential outlier. In particular, removing the two most extreme species by heart rate (the pygmy shrew at the high end and the elephant at the low end) does not change the qualitative conclusion. This analysis supports the robustness of the $p=0.09$ result and complements the Cook’s distance analysis in Extended Data Figure 4c, which shows all $D_{i}<4/n$.
Figure 9: Arrhenius correction: before and after, per species. Log-log scatter of maximum lifespan $L$ against heart rate $f_{H}$ for the 17 reptilian species in the dataset, showing both the uncorrected raw positions (crimson diamonds) and the Arrhenius-corrected positions (orange diamonds) after applying equation (71) with $E_{a}=0.65$ eV and $T_{\rm ref}=310$ K. Grey arrows connect each species’ raw and corrected positions, showing the direction and magnitude of the correction: correction shifts points horizontally to higher effective heart rates (equivalent to normalising the metabolic clock to mammalian body temperature), pulling the distribution toward the non-primate placental OLS line (dashed). Species with lower field body temperatures (e.g. Sphenodon punctatus at 293 K) receive larger corrections than warm-active species (e.g. Varanus komodoensis at 303 K). The correction reduces but does not eliminate the gap between reptiles and the mammalian baseline: a residual deviation of $\sim$0.07 dex in mean $\ell$ remains after correction (Extended Data Figure 3b), indicating that factors beyond the Arrhenius thermal effect contribute to the ectotherm-endotherm difference.
Figure 10: Distribution of $\ell$ before and after Arrhenius correction. Histogram of the PBTE invariant $\ell=\log_{10}(f_{H}\cdot L\cdot 525{,}960)$ for the 17 reptilian species before correction (crimson, raw $\bar{\ell}=8.61$) and after Arrhenius correction to $T_{\rm ref}=310$ K (orange, corrected $\bar{\ell}=8.93$). The vertical dashed line marks the non-primate placental mean $\bar{\ell}_{0}=8.995$. The Arrhenius correction shifts the reptile distribution by +0.32+0.32 dex on average, reducing the gap from the mammalian baseline from $0.38$ dex to $0.07$ dex—a reduction of approximately 82%. The residual gap of $0.07$ dex is statistically non-significant ($p=0.42$, Welch $t$-test) and may reflect the imprecision of field body temperature estimates, genuine residual physiological differences between ectotherms and endotherms, or both. The corrected distribution is more tightly concentrated than the raw distribution, suggesting that body temperature heterogeneity accounts for a substantial fraction of the within-clade scatter in reptile $\ell$ values.
Figure 11: Sensitivity of the corrected reptile mean $\ell$ to the activation energy $E_{a}$. Mean Arrhenius-corrected $\bar{\ell}^{\rm corr}$ for the 17 reptilian species as a function of the assumed activation energy $E_{a}\in[0.40,0.90]$ eV, computed using equation (71) with the species-specific field body temperatures in Extended Data Table 7. The grey shaded band shows the non-primate placental mean $\pm 1$ s.d. ($8.995\pm 0.160$), representing the range of $E_{a}$ values for which the corrected reptile mean would be statistically indistinguishable from the mammalian baseline. The vertical red dotted line marks the consensus metabolic activation energy $E_{a}=0.65$ eV from Gillooly et al. [26], at which the corrected reptile mean is $\bar{\ell}^{\rm corr}=8.93$. The corrected mean enters the mammalian $\pm 1$ s.d. band only for $E_{a}\gtrsim 0.78$ eV; at the consensus value $E_{a}=0.65$ eV a residual gap of $0.07$ dex persists. The linear dependence of $\bar{\ell}^{\rm corr}$ on $E_{a}$ reflects the Arrhenius correction formula; the sensitivity is $\approx 0.56$ dex per eV of $E_{a}$ at the mean reptile temperature deficit $\langle 1/T_{\rm field}-1/T_{\rm ref}\rangle\approx 1.25\times 10^{-4}$ K-1.
Figure 12: Standardised residuals versus fitted values. Standardised residuals plotted against fitted values $\hat{y}_{i}=\hat{\beta}\,x_{i}+\hat{\alpha}$ from the primary OLS regression of $\log_{10}L$ on $\log_{10}f_{H}$ for the $n=43$ non-primate placentals with directly measured heart rates. The horizontal dashed line at zero is shown for reference. No systematic pattern—such as curvature, funnelling, or banding—is evident in the scatter, indicating that the assumptions of linearity and homoscedasticity are well satisfied. The Breusch–Pagan test for heteroscedasticity gives $\chi^{2}=2.3$, $p=0.31$, confirming that the residual variance does not depend systematically on the fitted values. The absence of any trend in this plot also confirms that the log-log transformation is appropriate for this dataset and that no residual nonlinearity requires modelling. The two points with standardised residuals $|e_{i}|>2$ (one above and one below) are not influential observations, as confirmed by their Cook’s distances below the $4/n$ threshold (Extended Data Figure 4c).
Figure 13: Normal quantile–quantile plot. Normal Q–Q plot of the OLS residuals from the primary regression of $\log_{10}L$ on $\log_{10}f_{H}$ for the $n=43$ non-primate placental species. Sample quantiles of the residuals (blue points) are plotted against the theoretical quantiles of the standard normal distribution; the red line shows the expected relationship under exact normality. Points follow the reference line closely throughout the full range, including both tails, with no evidence of heavy tails, skewness, or bimodality. The Shapiro–Wilk test gives $W=0.97$, $p=0.19$, failing to reject normality at any conventional significance level. Together with the residuals-versus-fitted plot (Extended Data Figure 4a), this confirms that the standard OLS inferential framework—including the $t$-distribution approximation used in the $F$-test against $\beta=-1$— is appropriate for this dataset without requiring non-parametric corrections.
Figure 14: Cook’s distances. Cook’s influence distance $D_{i}$ for each of the $n=43$ non-primate placental species in the primary OLS regression. Cook’s distance measures the aggregate change in all fitted values when observation $i$ is deleted; large values indicate influential observations that exert disproportionate leverage on the regression coefficients. The red dashed horizontal line marks the conventional threshold $4/n=0.093$. All 43 observations satisfy $D_{i}<4/n$, with a maximum of $D_{\rm max}=0.079$. The three species with the highest Cook’s distances are labelled (indices 2, 13, and 36 in the dataset ordering of Extended Data Table 1). Removing all three simultaneously changes the OLS slope by less than 0.02, confirming that the regression result is not driven by influential outliers. This diagnostic also validates the leave-one-out analysis (Extended Data Figure 2c): the absence of high-$D_{i}$ points ensures that all leave-one-out $p$-values reflect genuine distributional properties of the dataset rather than the removal of pivotal observations.
Figure 15: Leverage versus standardised residuals. Standardised residuals plotted against leverage values $h_{ii}$ (diagonal elements of the hat matrix $\mathbf{H}=\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{-1}\mathbf{X}^{\top}$) for the $n=43$ non-primate placental species. The vertical dashed red line marks the high-leverage threshold $h=2p/n=0.093$, where $p=2$ (intercept and slope). No observation exceeds this threshold; the maximum observed leverage is $h_{ii}=0.148$ for a single species at the extreme of the heart-rate range. Although this one point has leverage above $2p/n$, its standardised residual is small ($|e_{i}|<1.5$) and its Cook’s distance is below $4/n$ (Extended Data Figure 4c), indicating that it is a high-leverage but non-influential observation. The broad spread of points across the residual range with no systematic dependence on leverage confirms the absence of leverage-residual confounding, a necessary condition for reliable OLS inference in this dataset.
@@ A.9. Assumptions used in the derivation
Heart rate type (see $f_{H}$ column header): directly measured resting values are used for all non-primate placentals, primates, marsupials, and birds. Three NP placental species with no published resting measurement (Rhinoceros unicornis, Dugong dugon, Orycteropus afer) have heart rates imputed from the allometric relation $f_{H}=241\,M^{-0.25}$ bpm [5] and are flagged with . For bats and cetaceans, $f_{H}^{\rm eff}$ is the duty-cycle-corrected time-average (see Sections 6.2 and 7.4). For ectotherms, $f_{H}^{\rm eff}$ is the Arrhenius-corrected value (see Section 9.3).